# Multimodal Generative Flows for LHC Jets

### v2

@@ Appendix C Architecture
To effectively model the multimodal nature of the data, we embed independently each input mode, $x^{d}_{t}$ and $k^{d}_{t}$, into a high-dimensional vector space $\mathbb{R}^{h_{\tt emb}}$ using Multi-layer perceptrons (MLP) consisting of two linear layers with a GELU non-linear activation function in between. To process the flavor tokens $k^{d}_{t}$ we replace the first layer of the MLP with a learnable lookup table implemented in PyTorch via nn.Embedding. The time variable $t\in[0,1]$ is encoded into $t_{\rm emb}\in\mathbb{R}^{h_{\tt emb}}$ using a Fourier feature (FF) embedding of Tancik et al. (2020). The resulting embeddings for each mode are then summed with the embedded time vector $t_{\rm emb}={\tt FF}(t)$ to form hidden kinematic and flavor representations:
@@
Given that the dataset consists of particle clouds, where the ordering of constituents is unimportant, it is essential to approximate the generators of the dynamical process with a permutation-equivariant architecture. In this work we employ a particle transformer architecture, or ParticleFormer for short, as the core component of our multimodal encoder. The ParticleFormer processes the embedded time variable along with the embedded particle-level features with a stack of multi-head self-attention blocks. Details are shown in Fig. 3 (a). We use a generic GPT-style self-attention block without causal masking and positional encoding to guarantee permutation equivariance. Particle transformers of this sort were first used for jet tagging on simulated data by Qu et al. (2022), producing state-of-the-art results when compared to previous methods such as GANs and graph neural networks.
@@ Appendix D Sampling algorithm details
As discussed in Sec. 2, the training dataset exhibits a pronounced class imbalance, with jets containing much more photons and charged hadrons compared to neutral hadrons and leptons. Such imbalances are well known to hinder classification performance [Johnson & Khoshgoftaar (2019)]. A common way to alleviate their impact is to recalibrate the posterior probabilities (61) through temperature scaling [Guo et al. (2017)]. Specifically, we introduce a temperature hyperparameter $T$ as
@@ Appendix E Experiment details
We train the baseline on the same target AOJ datasets. However, since flow-matching can only handle continuous variables, the flavor token of each particle is one-hot encoded into unit vectors in $\mathbb{R}^{S}$ representing flavor assignment probabilities. For the target jets, these probabilities are concatenated with the particle kinematics, forming an augmented continuous feature vector $x_{1}^{d}\in\mathbb{R}^{3+S}$. We generate source point-clouds by drawing each point from a standard Gaussian over $\mathbb{R}^{3+S}$. After generation, to ensure each particle has a unique flavor assignment, we apply an $\tt argmax$ operation to the generated assignment probabilities and tokenize back to $\mathcal{F}$. This strategy has been successfully employed in previous works [Birk et al. (2023); Araz et al. (2024)] for jet and event datasets. For our experiments, we implement the EPiC-FM encoder described in Buhmann et al. (2023a) depicted in Fig. 3 (b) with the following setup: $n_{\rm layers}=16$ EPiC layers with $h_{\rm loc}=256$ and $h_{\rm glob}=16$ for the local and global hidden dimensions. The resulting model has around 5.9 million parameters. Optimization is perform with the Adam algorithm [Kingma & Ba (2014)] with an effective batch size of 256 jets for a maximum of 1500 epochs. A cosine-annealing learning rate schedule is applied for the first 1000 epochs, decaying from $5\times 10^{-4}$ to $10^{-5}$, followed by 500 epochs at a fixed learning rate of $10^{-5}$.
@@
We set the gaussian smearing hyperparameter of the flow-matching component to $\sigma=10^{-5}$ and use a constant stochasticity parameter with $\beta=0.075$ for the multivariate telegraph process. During training, the time parameter $t$ is sampled uniformly from a slightly reduced unit interval $\epsilon,1-\epsilon]$ with $\epsilon=10^{-5}$, to prevent numerical instabilities caused at the time boundaries $t=0,1$ for the MJB model. We parameterize the combined generators $u_{t}^{\theta}\otimes\pi_{t}^{\theta}$ using the multimodal architecture introduced in Sec. 3 with a mid-fusion setup $(L_{1},L_{2},L)=(5,5,6)$. This configuration balances the intra-modal and cross-modal correlations between kinematics and flavor tokens in separate sub-modules with similar sizes. We fix the number of attention heads and hidden dimensions to $n_{\rm heads}=4$, $n_{\rm embd}=256$, and $n_{\rm inner}=512$, resulting in a model with approximately 5.6 million trainable parameters. Training is performed with the uncertainty weighted loss of Eq. (8). We use an uncertainty network with a 128-dimensional hidden state. Optimization is perform with the Adam algorithm [Kingma & Ba (2014)] with an effective batch size of 256 jets for a maximum of 1500 epochs. A cosine-annealing learning rate schedule is applied for the first 1000 epochs, decaying from $5\times 10^{-4}$ to $10^{-5}$, followed by 500 epochs at a fixed learning rate of $10^{-5}$.

### v3

@@ Appendix C Architecture
To effectively model the multimodal nature of the data, we embed independently each input mode, $x^{d}_{t}$ and $k^{d}_{t}$, into a high-dimensional vector space $\mathbb{R}^{h_{\tt emb}}$ using Multi-layer perceptrons (MLP) consisting of two linear layers with a GELU non-linear activation function in between. To process the flavor tokens $k^{d}_{t}$ we replace the first layer of the MLP with a learnable lookup table implemented in PyTorch via nn.Embedding. The time variable $t\in[0,1]$ is encoded into $t_{\rm emb}\in\mathbb{R}^{h_{\tt emb}}$ using a Fourier feature (FF) embedding of tancik2020fourier. The resulting embeddings for each mode are then summed with the embedded time vector $t_{\rm emb}={\tt FF}(t)$ to form hidden kinematic and flavor representations:
@@
Given that the dataset consists of particle clouds, where the ordering of constituents is unimportant, it is essential to approximate the generators of the dynamical process with a permutation-equivariant architecture. In this work we employ a particle transformer architecture, or ParticleFormer for short, as the core component of our multimodal encoder. The ParticleFormer processes the embedded time variable along with the embedded particle-level features with a stack of multi-head self-attention blocks. Details are shown in Fig. 4 (a). We use a generic GPT-style self-attention block without causal masking and positional encoding to guarantee permutation equivariance. Particle transformers of this sort were first used for jet tagging on simulated data by Qu:2022mxj, producing state-of-the-art results when compared to previous methods such as GANs and graph neural networks.
@@ Appendix D Sampling algorithm details
As discussed in Sec. 2, the training dataset exhibits a pronounced class imbalance, with jets containing much more photons and charged hadrons compared to neutral hadrons and leptons. Such imbalances are well known to hinder classification performance [johnson2019survey]. A common way to alleviate their impact is to recalibrate the posterior probabilities (63) through temperature scaling [guo2017calibration]. Specifically, we introduce a temperature hyperparameter $T$ as
@@ Appendix E Experiment details
We train the baseline on the same target AOJ datasets. However, since flow-matching can only handle continuous variables, the flavor token of each particle is one-hot encoded into unit vectors in $\mathbb{R}^{S}$ representing flavor assignment probabilities. For the target jets, these probabilities are concatenated with the particle kinematics, forming an augmented continuous feature vector $x_{1}^{d}\in\mathbb{R}^{3+S}$. We generate source point-clouds by drawing each point from a standard Gaussian over $\mathbb{R}^{3+S}$. After generation, to ensure each particle has a unique flavor assignment, we apply an $\tt argmax$ operation to the generated assignment probabilities and tokenize back to $\mathcal{F}$. This strategy has been successfully employed in previous works [Birk:2023efj; Araz:2024bom] for jet and event datasets. For our experiments, we implement the EPiC-FM encoder described in Buhmann:2023zgc depicted in Fig. 4 (b) with the following setup: $n_{\rm layers}=16$ EPiC layers with $h_{\rm loc}=256$ and $h_{\rm glob}=16$ for the local and global hidden dimensions. The resulting model has around 5.9 million parameters. Optimization is perform with the Adam algorithm [kingma2014adam] with an effective batch size of 256 jets for a maximum of 1500 epochs. A cosine-annealing learning rate schedule is applied for the first 1000 epochs, decaying from $5\times 10^{-4}$ to $10^{-5}$, followed by 500 epochs at a fixed learning rate of $10^{-5}$.
@@
We set the gaussian smearing hyperparameter of the flow-matching component to $\sigma=10^{-5}$ and use a constant stochasticity parameter with $\beta=0.075$ for the multivariate telegraph process. During training, the time parameter $t$ is sampled uniformly from a slightly reduced unit interval $\epsilon,1-\epsilon]$ with $\epsilon=10^{-5}$, to prevent numerical instabilities caused at the time boundaries $t=0,1$ for the MJB model. We parameterize the combined generators $u_{t}^{\theta}\otimes\pi_{t}^{\theta}$ using the multimodal architecture introduced in Sec. 3 with a mid-fusion setup $(L_{1},L_{2},L)=(5,5,6)$. This configuration balances the intra-modal and cross-modal correlations between kinematics and flavor tokens in separate sub-modules with similar sizes. We fix the number of attention heads and hidden dimensions to $n_{\rm heads}=4$, $n_{\rm embd}=256$, and $n_{\rm inner}=512$, resulting in a model with approximately 5.6 million trainable parameters. Training is performed with the uncertainty weighted loss of Eq. (9). We use an uncertainty network with a 128-dimensional hidden state. Optimization is perform with the Adam algorithm [kingma2014adam] with an effective batch size of 256 jets for a maximum of 1500 epochs. A cosine-annealing learning rate schedule is applied for the first 1000 epochs, decaying from $5\times 10^{-4}$ to $10^{-5}$, followed by 500 epochs at a fixed learning rate of $10^{-5}$.